A rock is thrown downward at an angle of -25 degrees (with respect to horizontal) off of a 35 m high cliff at a speed of 10 m/s. How far from the base of the cliff will the rock land? (The solution is 20.6 m, but I am not sure how to arrive at this answer.)
Answer #1

Vacolo I vz lomls Voy 35m old Voly = -losinas= – 4.226 m/s By motion and equation sautt hat² -35 = -4.2266-4.9t² It= 2. 276 s